# Equivariant plug-and-play image reconstruction

### v1

@@ 2 Related Work
A significant amount of work has recently been proposed to provide stable PnP algorithms. A popular line of research in that direction consists in constraining the denoiser’s architecture to ensure its stability when plugged into a PnP scheme. Some approaches propose to regularize the denoiser’s training loss with a term penalizing the Lipschitz constant of the denoiser [34, 29], which can be combined with architectural constraints [40, 17]. Other works propose to alter the optimization algorithm itself in order to ensure the its stability [27, 10, 48, 18]. In turn, ensuring convergence of the algorithm allows to ensure better transferability to new imaging tasks [41], but also to perform iteration-intensive tasks, such as sampling from the posterior distribution [23]. However, all the aforementionned methods come at the cost of either algorithmic modifications or strong constraints on the design of the denoiser, as opposed to the original, desnoiser agnostic approaches [43, 32].
@@
Equivariance to certain transformations, such as rotations or translation, has often been a desired property when designing handcrafted variational priors [36, 13, 35]. Geometric ensembling techniques have been known in the computer vision literature, where it has been shown to improve reconstruction quality for image super-resolution [42]. Yet, to the notable exception of [48, 41], this strategy went unnoticed in the PnP literature, and to the best of our knowledge, the role of equivariance has not been explored in the context of algorithms relying on implicit denoising priors. In a different line of work, but still in the context of imaging inverse problems, recent works have exploited equivariance for the design of unrolled network architectures [5], or the construction of self-supervised learning losses [7] (e.g., see the recent review [8]).
@@
More generally, the equivariance assumption appears as a natural assumption in a large number of application, and there exists a vast literature on building equivariant neural network architectures [11, 45, 4]. Typical applications involve segmentation on spherical manifolds, robotics, point cloud analysis, data augmentation to name a few [15, 7, 28, 33, 44, 51]. However, these networks often fail to perform as well as other state-of-the-art architectures. Moreover, it can be challenging to incorporate complex layers (e.g., upsampling/downsampling, attention-like layers, etc.) without breaking the equivariance of the resulting network. In this work, we provide a simple method for rendering any denoiser equivariant, without any architectural constraints.
@@ 3 PnP algorithms
Traditional variational approaches for solving (1) consists in reformulating it as a minimization problem. Following a maximum-a-posteriori approach, one can derive an estimate $\widehat{x}$ as
@@
$x_{k+1}=x_{k}-\gamma A^{\top}(Ax_{k}-y)-\gamma\lambda(x_{k}-\operatorname{D}(x_{k}))+\sqrt{2\gamma}\epsilon_{k},$ (ULA)
@@
While there exist many different variants of PnP, RED, and ULA, we here focus on their most standard formulations. Interestingly, these algorithms have shown impressive performance on a wide variety of imaging tasks while relying on Gaussian denoisers agnostic to the imaging modality of interest. Yet, these algorithms suffer from a lack of stability and potential divergence effects, hurting their versatility.
@@ 3.1 Proposed equivariant approach
Intuitively, imaging priors should have some invariance properties with respect to certain groups of transformations, such as rotations, translations, and reflections. We denote these transformations associated with a group $\mathcal{G}$, $\{T_{g}\}_{g\in\mathcal{G}}$ where $T_{g}\in\mathbb{R}^{n\times n}$ is a unitary matrix [^1] While it is possible to define group actions with non-unitary matrices [37], here we focus on the unitary matrices, which is the case of translations, rotations and reflections of images.. We say that $\operatorname{D}$ is equivariant to the group action $\{T_{g}\}_{g\in\mathcal{G}}$ if $\operatorname{D}(T_{g}x)=T_{g}\operatorname{D}(x)$ for all $x$ and $g\in\mathcal{G}$. At the algorithmic level, this requirement translates into the equivariance of the denoiser with respect to the transforms of interest: if $r(x)$ is a $\mathcal{G}$-invariant function, its proximal operator and gradient (if they exist) are necessarily $\mathcal{G}$-equivariant functions [5].
@@
The equivariant counterpart of the PnP, RED and ULA algorithms is thus simply obtained by replacing the denoiser $\operatorname{D}$ by a sample of the Monte Carlo estimate (4) at each iterate of the algorithm. Explicit versions of these algorithms can be found in the Supplementary Material (SM).
@@ 4 Theoretical analysis
We first show that if the signal distribution is $\mathcal{G}$-invariant, then for any denoiser $\operatorname{D}$, its averaged version 3 obtains an equal or better denoising performance. This can be shown by computing the expected $\ell_{2}$ error with respect to the signal and noise distributions, i.e.,
@@
is smaller than 1. If the spectra of the matrices $A^{\top}A$ and $\operatorname{D}$ are incoherent, i.e., $A^{\top}A$ and the Jacobian of the denoiser are diagonalized in different bases, the Lipschitz constant of 5 is likely to be smaller than that of $\operatorname{D}$. A similar stability phenomenon happens for the RED and ULA iterates due to the incoherence between the spectrum of the forward operator and the one of the equivariant denoiser. We formalize this intuition in the following proposition:
@@
Table 3 shows the local Lipschitz constant of the mapping in 5 for various popular denoisers, where $A$ is a random inpainting operator, which is not equivariant to shifts, rotations nor reflections. The constants are smaller than those shown in Table 2 and are below 1, ensuring contraction of the PnP iterates for that specific operator. The equivariant denoisers show smaller constants than the non-equivariant counterparts.
@@
Table 3: Local Lipschitz constant of PnP iteration $\|J_{x}(I-A^{\top}A)\|$, with $A$ being a random inpainting operator. Equivariant denoisers are obtained by averaging over the group of 90 rotations and reflections.
@@ 4.1 Non-linear example
Figure 2 illustrates the behaviour of the PnP sequence with and without the group averaging in 3 for two specific choices of $A$ in a 2D toy example, where $\mathcal{G}$ is the group of flips (see the SM for more details). In both cases, the algorithm involving the equivariant denoiser converges to a point close to the global minima associated with the prior $\lambda r(x)=\lambda\|B_{1}x\|_{1}$, whereas sequence generated by the non-equivariant algorithm diverges.
5 Experimental results
In this section, we evaluate the influence of the proposed equivariant approach for different algorithms and linear inverse imaging problems.
5.1 Problems considered
Image deblurring and image super-resolution
In this setting, we set $y=h*x+\epsilon$ in (1) where $h$ is a convolutional kernel and *** the circular convolution. We consider either Gaussian deblurring, in which case $h$ is a Gaussian kernel of standard deviation 1, and motion deblurring, in which case $h$ is the first kernel from [24]. In both cases, $\epsilon$ is a Gaussian noise with standard deviation $0.01$. In the image super-resolution (SR) setting, (1) writes as $y=\left(h*x\right)_{\Downarrow S}+\epsilon$ where $h$ is a Gaussian kernel of standard deviation 1, and $S$ denotes the undersampling factor. When $S=2$ (resp. $S=4$), $\epsilon$ is a Gaussian noise with standard deviation $0.01$ (resp. $0.05$). We test the proposed method on the Set3C dataset as well as the BSD10 dataset, a subset of 10 images from the BSD68 dataset.
MRI
Algorithms and backbone denoisers
We consider several backbone pretrained denoiser, namely DRUNet [48], SCUNet [49], SwinIR [25], DiffUNet [16] and DnCNN [47] as well as its Lipschitz version (LipDnCNN) from [29]. These architectures are representative of state-of-the-art image reconstruction architectures, involving both convolutional and attention layers. We stress that none of these networks is equivariant to translations or rotations, to the exception of DnCNN that show approximate translation equivariance. The choice of the backbone denoiser may influence the chosen algorithm. For instance, the DnCNN and SwinIR denoisers on which this article relies are trained for a fixed level of noise, limiting the ability to finetune the algorithm. Moreover, DiffUNet and SCUNet architectures can only be applied to color images, preventing their usage on the MRI problem. Unless mentioned otherwise, each algorithm is run for $10^{4}$ iterations.
Baselines
In this paper, we mainly investigate the influence of equivariant priors on the stability of PnP algorithms. We thus compare the proposed approach with variational approaches which can be seen as a class of convergent PnP algorithms. In particular, we use wavelet denoisers (i.e. $\operatorname{D}(x)=\operatorname{prox}_{\lambda\|\Psi\cdot\|_{1}}(x)$ for $\Psi$ a redundant wavelet dictionary), as well as total generalized variation (TGV) denoisers [3]. The PnP algorithm with Lipschitz denoiser corresponds to the method from [29] which ensures convergence of the PnP algorithm. We also compare our approach with the state-of-the-art DPIR algorithm [48] which runs a small number of steps of a half-quadratic splitting algorithm with fine-tuned decaying stepsizes. As a consequence, DPIR can be seen as representative of non-convergent, fine-tuned PnP algorithm.
@@ 5.1 Problems considered
Table 4: Mean reconstruction PSNR on different problems for the PnP algorithm with different backbone denoisers. Three top rows show results with non-equivariant denoisers, while the three middle rows show the proposed equivariant modification. The last three rows provide standard reconstruction baselines. “div.” indicated divergence of the method.
@@ 5.2 Stability of PnP and RED algorithms
Figure 4: Average PSNR (left) and convergence criterion $\|x_{k+1}-x_{k}\|/\|x_{k}\|$ (right) for 3 different imaging problems as a function of the iteration number with different backbone denoisers plugged in the PnP algorithm. Top row: DnCNN, middle row: DRUNet, bottom row: Lipschitz DnCNN.
Figure 4 shows the PSNR and convergence criterion $\|x_{k+1}-x_{k}\|/\|x_{k}\|$ on different problems and backbone architecture. For each of these problems, the denoising level $\sigma$ in (PnP) is set to $\sigma=0.01$ (resp. $\sigma=0.015$) for DnCNN (resp. DRUNet) backbone architectures. We notice that the (PnP) algorithm with equivariant denoiser shows a more stable PSNR along iterations than its non-equivariant counterpart. In particular, we observe in the right panel of Figure 4 that the convergence rate of the (PnP) algorithm with $\mathcal{G}$-equivariant DnCNN matches the behaviour of the algorithm with Lipschitz denoisers. Lastly, we notice that the proposed equivariant approach also benefits convergent (PnP) algorithms relying on Lipschitz backbone denoisers. Reconstruction metrics are provided in Table 4.
Groundtruth $A^{\top}y$ TGV (29.31) Wavelets (29.99)
Refer to caption Refer to caption Refer to caption Refer to caption
DnCNN (22.34) SCUNet (12.56) DRUNet (8.93) SwinIR (27.83)
Standard Refer to caption Refer to caption Refer to caption Refer to caption
DnCNN (30.52) SCUNet (13.89) DRUNet (29.44) SwinIR (27.82)
Equiv. Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Gaussian deblurring on a BSD10 sample (detail) for different denoising backbone plugged in the PnP algorithm. Middle row: standard algorithms; bottom row: equivariant algorithms.
In the case of the (RED) algorithm, we observe similar behaviours. Reconstructions on a $\times$2 SR problem are shown in Figure 6 with associated convergence plots. In the case of a non-equivariant DRUNet backbone, the reconstructed image shows important geometric artifacts that disappear in the equivariant case.
@@
DiffUNet DRUNet Refer to caption Refer to caption Refer to caption Observed Standard (21.55) Standard (20.75) Refer to caption Refer to caption Refer to caption Groundtruth (PSNR) Equivariant (22.26) Equivariant (24.94)
@@
Figure 6: Left: Results of the RED algorithm on a ×$\times 2$ SR problem for different backbone denoisers. Middle column: DiffUNet; right column: DRUNet. Top row: standard algorithm; Bottom row: equivariant algorithm. Right: Associated convergence criterion.
@@ 5.3 Interplay with the kernel of $A$
In the previous section, we have seen that equivariant denoisers can prevent the emergence of artifacts along the iteration of (PnP). Despite their unnatural aspect, these artifacts are not incompatible with good data-fidelity measures. In fact, PnP algorithm offers no control over $\ker(A)$, which is nontrivial by nature of the ill-posed inverse problem (1); artifacts appearing during the reconstruction are therefore likely to belong to $\ker(A)$.
@@
This phenomenon can be illustrated in the case of MRI, where $\ker(A)$ corresponds to the non-sampled subspace of the Fourier domain. This is illustrated in Figure 7 showing the reconstruction with and without the proposed equivariant algorithmic update. In the standard (non-equivariant) setting, mild artifacts appear between iteration $j=10^{4}$ and $j=10^{3}$. These may appear unnoticed in $x_{j}$; however, they clearly appear when plotting the difference $x_{j}-x_{i}$. Interestingly, the Fourier spectrum of these artifacts shows significantly more energy at frequencies that were not sampled. The same experiment with the equivariant algorithm shows no such artifacts and a more uniform Fourier spectrum.

### v2

@@ 2 Related Work
A significant amount of work has recently been proposed to provide stable PnP algorithms. A popular line of research in that direction consists in constraining the denoiser’s architecture to ensure its stability when plugged into a PnP scheme. Some approaches propose to regularize the denoiser’s training loss with a term penalizing the Lipschitz constant of the denoiser [34, 29], which can be combined with architectural constraints [40, 17]. Other works propose to alter the optimization algorithm itself in order to ensure its stability [27, 10, 48, 18]. In turn, ensuring convergence of the algorithm allows to ensure better transferability to new imaging tasks [41], but also to perform iteration-intensive tasks, such as sampling from the posterior distribution [23]. However, all the aforementionned methods come at the cost of either algorithmic modifications or strong constraints on the design of the denoiser, as opposed to the original, denoiser agnostic approaches [43, 32].
@@
Equivariance to certain transformations, such as rotations or translations, has often been a desired property when designing handcrafted variational priors [36, 13, 35]. Geometric ensembling techniques have been known in the computer vision literature, where it has been shown to improve reconstruction quality for image super-resolution [42]. In the case of linear denoisers, symmetrization of the Jacobian has been shown to improve performance of linear denoisers [milanfar2013symmetrizing]. Yet, to the notable exception of [48, 41], this strategy went unnoticed in the PnP literature, and to the best of our knowledge, the role of equivariance has not been explored in the context of algorithms relying on implicit denoising priors. In a different line of work, but still in the context of imaging inverse problems, recent works have exploited equivariance for the design of unrolled network architectures [5], or the construction of self-supervised learning losses [7] (e.g., see the recent review [8]).
@@
More generally, invariance of the underlying prior appears as a natural assumption in a large number of applications, and there exists a vast literature on building equivariant neural network architectures [11, 45, 4]. Typical applications involve segmentation on spherical manifolds, robotics, point cloud analysis, data augmentation to name a few [15, 7, 28, 33, 44, 51]. However, these networks often fail to perform as well as other state-of-the-art architectures. Moreover, it can be challenging to incorporate complex layers (e.g., upsampling/downsampling, attention-like layers, etc.) without breaking the equivariance of the resulting network. In this work, we provide a simple method for rendering any denoiser equivariant, without any architectural constraints.
@@ 3 PnP algorithms
Traditional variational approaches for solving (1) consist in reformulating it as a minimization problem. Following a maximum-a-posteriori approach, one can derive an estimate $\widehat{x}$ as
@@
$x_{k+1}=x_{k}-\gamma A^{\top}(Ax_{k}-y)-\gamma\lambda(x_{k}-\operatorname{D}(x_{k}))\\ +\sqrt{2\gamma}\epsilon_{k},$ (3)
@@
While there exist many different variants of PnP, RED, and 3, we here focus on their most standard formulations. Interestingly, these algorithms have shown impressive performance on a wide variety of imaging tasks while relying on Gaussian denoisers agnostic to the imaging modality of interest. Yet, these algorithms suffer from a lack of stability and potential divergence effects, hurting their versatility.
@@ 3.1 Proposed equivariant approach
Intuitively, imaging priors should have some invariance properties with respect to certain groups of transformations, such as rotations, translations, and reflections. We denote transformations associated with a group $\mathcal{G}$ as $\{T_{g}\}_{g\in\mathcal{G}}$ where $T_{g}\in\mathbb{R}^{n\times n}$ is a unitary matrix [^1] While it is possible to define group actions with non-unitary matrices [37], here we focus on the unitary matrices, which is the case of translations, rotations and reflections of images.. We say that $\operatorname{D}$ is equivariant to the group action $\{T_{g}\}_{g\in\mathcal{G}}$ if $\operatorname{D}(T_{g}x)=T_{g}\operatorname{D}(x)$ for all $x$ and $g\in\mathcal{G}$. At the algorithmic level, this requirement translates into the equivariance of the denoiser with respect to the transforms of interest: if $r(x)$ is a $\mathcal{G}$-invariant function, its proximal operator and gradient (if they exist) are necessarily $\mathcal{G}$-equivariant functions [5].
@@
The equivariant counterpart of the PnP, RED and 3 algorithms is thus simply obtained by replacing the denoiser $\operatorname{D}$ by a sample of the Monte Carlo estimate (5) at each iterate of the algorithm. Explicit versions of these algorithms can be found in the Supplementary Material (SM).
@@ 4 Theoretical analysis
We first show that if the signal distribution is $\mathcal{G}$-invariant, then for any denoiser $\operatorname{D}$, its averaged version 4 obtains an equal or better denoising performance. This can be shown by computing the expected $\ell_{2}$ error with respect to the signal and noise distributions, i.e.,
@@
is smaller than 1. If the spectra of the matrices $A^{\top}A$ and $\operatorname{D}$ are incoherent, i.e., $A^{\top}A$ and the Jacobian of the denoiser are diagonalized in different bases, the Lipschitz constant of 6 is likely to be smaller than that of $\operatorname{D}$. A similar stability phenomenon happens for the RED and 3 iterates due to the incoherence between the spectrum of the forward operator and the one of the equivariant denoiser. The following proposition formalizes this intuition in the case of linear denoisers:
@@
Table 3 shows the local Lipschitz constant of the mapping in 6 for various popular non-linear denoisers, where $A$ is a random inpainting operator, which is not equivariant to shifts, rotations nor reflections. The constants are smaller than those shown in Table 2 and are below 1, ensuring contraction of the PnP iterates for that specific operator. The equivariant denoisers show smaller constants than the non-equivariant counterparts.
@@
Table 3: Local Lipschitz constant of PnP iteration $\|J_{x}(I-A^{\top}A)\|$, with $A$ a random inpainting operator. Equivariant denoisers are obtained by averaging over the group of 90 rotations and reflections.
@@ 4.1 Non-linear example
Figure 2 illustrates the behaviour of the PnP sequence with and without the group averaging in 4 for two specific choices of $A$ in a 2D toy example, where $\mathcal{G}$ is the group of flips (see the SM for more details). In both cases, the algorithm involving the equivariant denoiser converges to a point close to the global minima associated with the prior $\lambda r(x)=\lambda\|B_{1}x\|_{1}$, whereas sequence generated by the non-equivariant algorithm diverges.
@@
Table 4: Mean reconstruction PSNR for various image restoration problems using the PnP algorithm with different backbone denoisers. The first, third, and fifth rows present results for non-equivariant denoisers (LipDnCNN, DnCNN, and DRUNet, respectively), while their equivariant counterparts (Eq. LipDnCNN, Eq. DnCNN, and Eq. DRUNet) are shown in the second, fourth, and sixth rows. The bottom four rows offer benchmarks with standard reconstruction methods. The notation ”div.” indicates cases where the method diverged.
5 Experimental results
In this section, we evaluate the influence of the proposed equivariant approach for different algorithms and linear inverse imaging problems. Our implementation333available at https://github.com/matthieutrs/EquivariantPnP. relies on the DeepInverse library [deepinv].
5.1 Problems considered
Image deblurring and image super-resolution
In this setting, we set $y=h*x+\epsilon$ in (1) where $h$ is a convolutional kernel and * the circular convolution. We consider either Gaussian deblurring, in which case $h$ is a Gaussian kernel of standard deviation 1, and motion deblurring, in which case $h$ is the first kernel from [24]. Unless specified otherwise, $\epsilon$ is a Gaussian noise with standard deviation $0.01$. In the image super-resolution (SR) setting, (1) writes as $y=\left(h*x\right)_{\Downarrow S}+\epsilon$ where $h$ is a Gaussian kernel of standard deviation 1, and $S$ denotes the undersampling factor. When $S=2$ (resp. $S=4$), $\epsilon$ is a Gaussian noise with standard deviation $0.01$ (resp. $0.05$). We test the proposed method on the Set3C dataset as well as the BSD10 dataset, a subset of 10 images from the BSD68 dataset.
MRI
Algorithms and backbone denoisers
We consider several backbone pretrained denoiser, namely DRUNet [48], SCUNet [49], SwinIR [25], DiffUNet [16], DnCNN [47] as well as its 1-Lipschitz version (LipDnCNN) from [29] and the gradient-step denoiser GSNet [17]. These architectures are representative of state-of-the-art image reconstruction architectures, involving both convolutional and attention layers. We stress that none of these networks is equivariant to translations or rotations, to the exception of DnCNN that show approximate translation equivariance. The choice of the backbone denoiser may influence the chosen algorithm. For instance, the DnCNN and SwinIR denoisers on which this article relies are trained for a fixed level of noise, limiting the ability to finetune the algorithm. Moreover, DiffUNet and SCUNet architectures can only be applied to color images, preventing their usage on the MRI problem. Unless specified, each algorithm is run for $10^{4}$ iterations.
Baselines
In this paper, we mainly investigate the influence of equivariant priors on the stability of PnP algorithms. We thus compare the proposed approach with variational approaches which can be seen as a class of convergent PnP algorithms. In particular, we use wavelet denoisers (i.e. $\operatorname{D}(x)=\operatorname{prox}_{\lambda\|\Psi\cdot\|_{1}}(x)$ for $\Psi$ a redundant wavelet dictionary), as well as total generalized variation (TGV) denoisers [3]. The PnP algorithm with Lipschitz denoiser corresponds to the method from [29] which ensures convergence of the PnP algorithm. We also compare our approach with the state-of-the-art DPIR algorithm [48] which runs a small number of steps of a half-quadratic splitting algorithm with fine-tuned decaying stepsizes. As a consequence, DPIR can be seen as representative of non-convergent, fine-tuned PnP algorithm. We also compare our results with the gradient-step PnP algorithm [17], leveraging a nonconvex implicit prior and backtracking and that is representative of state-of-the-art convergent PnP algorithms.
5.2 Stability of PnP and RED algorithms
@@
Figure 4: Average PSNR (left) and convergence criterion $\|x_{k+1}-x_{k}\|/\|x_{k}\|$ (right) for 3 different imaging problems as a function of the iteration number with different backbone denoisers plugged in the PnP algorithm. Top row: DnCNN, middle row: DRUNet, bottom row: 1-Lipschitz DnCNN.
Figure 4 shows the PSNR and convergence criterion $\|x_{k+1}-x_{k}\|/\|x_{k}\|$ on different problems and backbone architecture. For each of these problems, the denoising level $\sigma$ in (PnP) is set to $\sigma=0.01$ (resp. $\sigma=0.015$) for DnCNN (resp. DRUNet) backbone architectures. We notice that the (PnP) algorithm with equivariant denoiser shows a more stable PSNR along iterations than its non-equivariant counterpart. In particular, we observe in the right panel of Figure 4 that the convergence rate of the (PnP) algorithm with $\mathcal{G}$-equivariant DnCNN matches the behaviour of the algorithm with Lipschitz denoisers. Lastly, we notice that the proposed equivariant approach also benefits convergent (PnP) algorithms relying on Lipschitz backbone denoisers. We further stress that reconstructions obtained with the equivariant Monte-Carlo estimates are consistent with those obtained with deterministic Reynolds averaging (see Tab. 5).
In the case of the (RED) algorithm, we observe similar behaviours. Reconstructions on a $\times$2 SR problem are shown in Figure 6 and associated convergence plots can be found in Figure 7. In the case of a non-equivariant DRUNet backbone, the reconstructed image shows important geometric artifacts that disappear in the equivariant case.
@@
Groundtruth
$y$
TGV (28.40)
Wavelets (28.64)
Refer to caption Refer to caption Refer to caption Refer to caption
SCUNet (23.54)
SwinIR (27.85)
DRUNet (9.58)
GSNet (29.22)
DnCNN (29.66)
Standard Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
SCUNet (25.14)
SwinIR (27.83)
DRUNet (29.22)
GSNet (29.25)
DnCNN (30.36)
Equiv. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Gaussian deblurring with standard deviation $\sigma=0.02$ on a BSD10 sample (detail) for different denoising backbone plugged in the (PnP) algorithm.
DiffUNet DRUNet
Refer to caption Refer to caption Refer to caption
Observed Standard (21.55) Standard (20.75)
Refer to caption Refer to caption Refer to caption
Groundtruth (PSNR) Equivariant (22.26) Equivariant (24.94)
Figure 6: Results of the RED algorithm on a ×$\times 2$ SR problem for different backbone denoisers. Middle column: DiffUNet; right column: DRUNet. Top row: standard algorithm; Bottom row: equivariant algorithm.
@@
Figure 7: Evolution of the PSNR along the RED algorithm for different backbone denoisers associated to the reconstructions shown in Figure 6.
@@ 5.3 Interplay with the kernel of $A$
In the previous section, we have seen that equivariant denoisers can prevent the emergence of artifacts along theiterations of (PnP). Despite their unnatural aspect, these artifacts are not incompatible with good data-fidelity measures. In fact, PnP algorithms offer no control over $\ker(A)$, which is nontrivial by nature of the ill-posed inverse problem (1); artifacts appearing during the reconstruction are therefore likely to belong to $\ker(A)$.
This phenomenon can be illustrated in the case of MRI, where $\ker(A)$ corresponds to the non-sampled subspace of the Fourier domain. In this case, $A$ is not rotation equivariant and Proposition 3 suggests that enforcing equivariance of the prior improves the stability. This is illustrated in Figure 8 showing the reconstruction with and without the proposed equivariant algorithmic update. In the standard (non-equivariant) setting, mild artifacts appear between iteration $i=10^{3}$ and $j=10^{4}$. These may appear unnoticed in $x_{j}$; however, they clearly appear when plotting the difference $x_{j}-x_{i}$. Interestingly, the Fourier spectrum of these artifacts shows significantly more energy at frequencies that were not sampled. The same experiment with the equivariant algorithm shows no such artifacts and a more uniform Fourier spectrum.